OpenStudy (anonymous):

PLEASE Can Anyone help me? Iv'e been waiting and need help ASAP!! Simplify: (5a)(6b)(2c). (Points : 1) 60(a+b+c) 13(a+b+c) 13abc 60abc Question 7. 7. Simplify: 4 6/7 + 9 1/4 + 2 3/4 + 1/7. (Points : 1) 19 18 17 Question 8. 8. Simplify: 5(2m + 6n - 3n) (Points : 1) 10m + 15n 10m + 6n - 3n 7m + 11n - 2n 10m + 11n - 2n Question 9. 9. Simplify: 12(5c - 3). (Points : 1) 17c - 9 60c - 3 60c + 36 60c - 36

OpenStudy (anonymous):

OK first off is this homework test or quiz?

OpenStudy (anonymous):

home work

OpenStudy (anonymous):

Oh ok well the first one is B second one is D and the thuird one i don't knmow

OpenStudy (anonymous):

that doesnt really help me

OpenStudy (kl0723):

@lee_actress you will need to distribute every number and then every letter... for # 1 you have 5*6*2=60 and a*b*c = abc all together makes 60abc :)

OpenStudy (anonymous):

okay so it's not what he said?

OpenStudy (kl0723):

not at all...

OpenStudy (anonymous):

thank you can you help me with the others?

OpenStudy (kl0723):

yes... does the question in the second problem look like the following? \[4\frac{ 7 }{ 6 }+9\frac{ 1 }{ 4}... etc???\]

OpenStudy (anonymous):

no it's exactly like this Simplify: 4 6/7 + 9 1/4 + 2 3/4 + 1/7. (Points : 1) 19 18 17

OpenStudy (anonymous):

but if the other method works for u go ahaed

OpenStudy (kl0723):

so that is \[\frac{ 46 }{ 7 }...\]

OpenStudy (anonymous):

as the answer?

OpenStudy (kl0723):

no.. the question.. it just looks like a mixed number

OpenStudy (anonymous):

let's skip this one can you help me with this one? Simplify: 5(2m + 6n - 3n) (Points : 1) 10m + 15n 10m + 6n - 3n 7m + 11n - 2n 10m + 11n - 2n

OpenStudy (anonymous):

how did you do the 46/7 like that

OpenStudy (kl0723):

how about you post a picture of the actual problem... can u do that?

OpenStudy (anonymous):

no i can only copy and paste

OpenStudy (kl0723):

ok I'll check # 2 now... one sec

OpenStudy (anonymous):

okay i have 10 more minutes

OpenStudy (kl0723):

ok... I got it, those are mixed numbers and need to be converted to improper fractions

OpenStudy (anonymous):

okay

OpenStudy (kl0723):

so solve it like this \[4\frac{ 7}{ 6}= (4*6)+7 = 31/6\]

OpenStudy (kl0723):

that means take 4 multiply it times the bottom number (6) and add top number (7) to that

OpenStudy (anonymous):

okay

OpenStudy (kl0723):

equals 31 and then divide it by the same bottom number :) 31/6

OpenStudy (anonymous):

thank you!! I hav3 minutes here is the last one!! Simplify: 5(2m + 6n - 3n) (Points : 1) 10m + 15n 10m + 6n - 3n 7m + 11n - 2n 10m + 11n - 2n

OpenStudy (kl0723):

I'll help with the answer since you have little time but now you know how to convert... #2 is option 3

OpenStudy (kl0723):

10m+30n-15n = 10m+15n so that is option 1

OpenStudy (kl0723):

last is 60c-36 option 4

OpenStudy (anonymous):

Thank you i have this last one!! Simplify: 5x + 2y - 3x + y. (Points : 1) 2x + 2y 2x +3y -15x + 2y -30xy

OpenStudy (kl0723):

add like terms x's with x's and y's with y's so that gives you 2x+3y

OpenStudy (kl0723):

option 2

OpenStudy (anonymous):

Thank you SO MUCH FOR ALLL YOUR HELP!!!! :)

OpenStudy (kl0723):

yw :)